OpenStudy (anonymous):

What is the solution to the equation 36^3c = sqr 6^(8c - 6) ?

sam (.sam.):

\[36^{3 c}=\sqrt{6}^{8 c-6}\]

OpenStudy (anonymous):

yes

sam (.sam.):

\[6^2=36 ~~~~~ \sqrt6 =6^{1/2}\] ---------------------- \[6^{6c}=6^{(1/2)(8c-6)}\] 6c=(1/2)(8c-6) Solve

OpenStudy (anonymous):

c=-4/3?

sam (.sam.):

no, \[6c=\frac{8c-6}{2}\]

OpenStudy (anonymous):

-3/2?

sam (.sam.):

yes :)

OpenStudy (anonymous):

THANKS!